A continuous random variable x has density function
f(x)= 3/4x(2-x) , 0<x<2
Find p(x>1/4)
1
Expert's answer
2020-09-07T19:54:02-0400

"f(x)=\frac{3x(2-x)}{4}" , "0<x<2"

Check if it is a pdf

"\int_0^2\frac{3x(2-x)}{4}dx"

"=[\frac{3x^2}{4}-\frac{x^3}{4}]_0^2"

"=(3-2)-0=1" (it is a pdf)

"P(x>\frac{1}{4})=\int_{\frac{1}{4}}^2\frac{3x(2-x)}{4}dx"

"=[\frac{3x^2}{4}-\frac{x^3}{4}]_{\frac{1}{4}}^2"

"=[3-2]-[\frac{1}{64}(3-\frac{1}{4})]"

"=1-\frac{11}{256}=\frac{245}{256}=0.957"